OpenStudy (anonymous):

In the 30°-60°-90° triangle above, y = 6. Find the other two lengths

OpenStudy (anonymous):

can u post the triangle

OpenStudy (anonymous):

yea hold one

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

\[\large \sf Hypotenuse~=~y\\\large Long~Leg~(LL)~=~z\\\large Short~Leg~(SL)~=~x\] \[\large \sf SL~=~\frac{1}{2} \times Hypotenuse\] \[\large \sf LL~=~SL \times \sqrt{3}\]

OpenStudy (anonymous):

thank you now i know how to solve it

OpenStudy (anonymous):

No problem :)